Errata for Surreal Numbers (last updated 09 Feb 2008)

Page 24, line 6 from the bottom or equal to or like Page 36, line 5 from the bottom rule (3) rule (2) Page 45, line 8 look’s looks Page 52, line 7 from the bottom ), ). Page 86, lines 14 and 15 II(X_{L}, y, z), II(x, Y_{R}, z). II(y, X_{L}, z), II(Y_{R}, x, z). Page 86, the bottom two lines IV(Z_{R}, z, x), IV(X_{L}, y, z), IV(x, Y_{R}, z). IV(Z_{R}, z, x). Page 99, line 2

(4) (3) Page 110, lines 9 and 10 equal to 2 like 2 Page 111, lines 4 and 5 ({x, x − x^{2} + x^{3}, x − x^{2} + x^{3} − x^{4} + x^{5}, … }, {x − x^{2}, x − x^{2} + x^{3} − x^{4}, … }) ({x − x^{2}, x − x^{2} + x^{3} − x^{4}, … }, {x, x − x^{2} + x^{3}, x − x^{2} + x^{3} − x^{4} + x^{5}, … }) Page 112, line 5 from the bottom The theory [See his incredible book On Numbers and Games (London: Academic Press, 1976), as well as the two-volume sequel by Berlekamp, Guy, and Conway, Winning Ways (London: Academic Press, 1982).] The theory Page 113, line 23 Grundlagen der Mathematik Grundlagen der Analysis Page 117, lines 3 and 4 of problem 18 X_{L} has a greatest element if and only if X_{R} has a least element. X_{L} has a greatest element or is null if and only if X_{R} has a least element or is null. Page 119, append a quotation following the text

Detection is, or ought to be, an exact science,

and should be treated in the same cold and unemotional manner.

You have attempted to tinge it with romanticism, which produces much the same effect as if you

worked a love-story or an elopement into the fifth proposition of Euclid.

— SHERLOCK HOLMES (Sign of The Four, Chapter 1)